Consider the following two alternatives: Year A 0 -$2,000 +$800 +$800 3 +$800 B $2,800 +$1,100 +$1,100 +$1,100 2 If the minimPlease, show the formulas and step by step on how to find the answer!!
Answer #1

Use PW analysis

PW of A = -2000 + 800(P/A, 5%, 3)

= -2000 + 800*(1.05^3-1)/(0.05*1.05^3)

= 178.60

PW of B = -2800 + 1100(P/A, 5%, 3)

= -2800 + 1100*(1.05^3-1)/(0.05*1.05^3)

= 195.57

Since PW of B is greater than PW of A select B.